OpenStudy (anonymous):

what if the calculation continued forever Calculate: 1+(1/1)+(1/1x2)+(1/1x2x3)+(1/1x2x3x4)+... +...+...

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no idea is there a method you are supposed to use?

OpenStudy (anonymous):

i believe its eulers formula

OpenStudy (anonymous):

\[1+\frac{1}{1!}+\frac{1}{2!}+\frac{1}{3!}+...\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh , it is \(e\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you replace \(1\) by \(x\) that is the expansion of \(e^x\)

OpenStudy (anonymous):

could you please show it to me

OpenStudy (anonymous):

\[e^x=\sum_{n=0}^{\infty}\frac{x^n}{n!}\]

OpenStudy (anonymous):

so \[e^1=\sum_{n=1}^{\infty}\frac{1}{n!}\]

OpenStudy (anonymous):

is that the formula?

OpenStudy (anonymous):

that is the power series expansion for \(e^x\) yes

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

that is a very very common and famous one easy to remember too, so memorize it

OpenStudy (anonymous):

so how can we find the answer?

OpenStudy (anonymous):

i still dont know how to compute it

OpenStudy (anonymous):

i mean how do you compute the part on your right hand side?